# Rhetorical XAI: Explaining AI’s Benefits as well as its Use via Rhetorical Design

### v1

@@ B.1. Reviewed articles
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Alqaraawi et al. (2020) Isolated presentation Logical: Subjective metrics Frequency of feature mentioned
Credible: Subjective metrics Self-confidence
Alvarado et al. (2022) Isolated presentation; Expressive interaction Credible: Qualitative feedback Increased sense of user control and awareness
Emotional: Qualitative feedback Increased sense of user enjoyment
Bansal et al. (2019a) Adaptive presentation Logical: Objective metrics User decision-making error boundary across trials
Bansal et al. (2019b) Adaptive presentation Logical: Objective metrics User decision-making error boundary across trials
Bansal et al. (2020) Adaptive presentation Logical: Objective metrics Human-AI agreement; Subjective metrics Explanation is useful, AI assistant is useful; Qualitative themes Used AI as a starting references point; Double-checked after they made own decisions
@@
Credible: Qualitative themes Explicitly acknowledged they used the confidence [score]; Was more likely to accept AI above the [confidence] threshold
Ben David et al. (2021) Isolated presentation; Protective framing Logical: Objective metrics Acceptance of model advice
Credible: Subjective metrics Willingness to pay
Emotional: Independent variables User algorithmic aversion
@@
Bernardo and Seva (2023) Isolated presentation; Explanation affect; Emotional cues Logical: Subjective metrics Perceived usefulness;
Credible: Subjective metrics Perceived trust, Reliance;
Emotional: Subjective metrics AI anxiety, Incidental emotion
Bove et al. (2022) Mechanistic presentation Logical: Subjective metrics User understanding of the system (e.g., local feature understanding, difference between local and global explanations)
Credible: Subjective metrics Explanation satisfaction
Cai et al. (2019a) Isolated presentation Logical: Subjective metrics System understanding, capability, and benevolence
Cai et al. (2019b) Mechanistic presentation; Explanation as retriever Logical: Subjective metrics Diagnostic utility, Future use, Overall preference; Qualitative themes Tracking the likelihood of a decision hypothesis, Generating new ideas; Reducing problem complexity
Credible: Subjective metrics Trust
B.2. Reviewed articles (continued)
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Cau and Spano (2025) Mechanistic presentation; Technical assurance Logical: Objective metrics Appropriate reliance; Subjective metrics User understanding and usefulness of explanations Qualitative themes Local explanations are fast and time-saving for decisions; Difficulty in understanding AI assigned weights on attributes, Overwhelmed by information
Credible: Subjective metrics User confidence, Trustworthiness of explanations; Qualitative themes Data-centric explanations were trustworthy with clear guidelines
Chen et al. (2023b) Isolated presentation Logical: Objective metrics Appropriate reliance; Subjective metrics Self-preferences; Qualitative themes Example-based explanations provide additional context and support inductive reasoning compared to feature-based; Example-based explanations led to less disruption of people intuition and provide strong signals of AI unreliability
Díaz-Rodríguez and Pisoni (2020) Authority framing; Cultural relatedness Credible: Qualitative themes Embedding expert knowledge for explanation
Emotional: Qualitative themes Communicating emotional depth in explanations
Ebermann et al. (2023) Isolated presentation; Protective framing Logical: Subjective metrics Evaluation of AI support
Credible Subjective metrics: Willingness to use AI
Emotional Subjective metrics: Mood; Independent variables: Cognitive misfit
Ehsan et al. (2021) Isolated presentation; Social endorsement Credible: Qualitative themes Increased social validation, transitive trust, and temporal relevance
Ehsan et al. (2024a) Isolated presentation; Numerical association Logical: Subjective metrics Perceived intelligence, Understandability
Credible: Subjective metrics Confidence, Friendliness, Second chance
Evans et al. (2022) Isolated presentation; Graphical synergy Logical: Subjective metrics Explanations are understandable, Explanations are useful; Qualitative themes Counterfactual explanations enable correct understanding of model behavior, Concept explanations seen as intuitive understandable but lack of precision; Saliency maps are visually appealing but risk promoting positive confirmation bias and obscure information.
Credible: Qualitative themes Confidence labels are intuitive cues for model reasoning
B.3. Reviewed articles (continued)
Paper XAI Goals Design Strategies Mapping User Effect to Rhetorical Appeals
How Why
Freire et al. (2023) Isolated presentation; Emotional cues Credible: Objective metrics Acceptance of AI advice
Emotional: Subjective metrics User engagement
@@ B.3. Reviewed articles (continued)
Gaole et al. (2025) Adaptive presentation; Mechanistic presentation Logical: Objective metrics Objective feature understanding, Appropriate reliance based on human-AI agreement; Subjective metrics Perceived explanation utility
@@
Credible: Subjective metrics User trust of the AI system, User confidence
Emotional: Subjective metrics: User engagement
Kenny et al. (2021b) Isolated presentation; Explanation as retriever Logical: Subjective metrics Perceived correctness, Reasonability
Emotional: Subjective metrics Trust, Satisfaction
Kim et al. (2024) Mechanistic presentation; Explanation as recommender Logical: Qualitative themes Increased usability for clinicians, Stronger desire to identify the types and causes of errors made by patients

### v2

@@ Appendix B Narrative Review Process and Corresponding Coding Scheme for Reviewed Articles
Table 4. Reviewed XAI Design Papers: Mapping Design Strategies to Rhetorical Appeals
Paper (N=41) Mapping User Effect to Rhetorical Appeals
Alqaraawi et al. (2020) How AI Works: Isolated presentation
Logical Participant mentions of saliency maps features to understand the system:
p280: “…The rationale for this was that we aimed to compare how frequently participants mentioned concepts related features that saliency maps could potentially highlight…Our results show that when saliency maps were shown, participants predicted the outcome of the classifier significantly more accurately…However, even with the presence of saliency maps, success rates were still relatively low (60.7%)” p282-283: “…Furthermore, it is hard to get a quantifiable measure of the importance of individual features in an image. Again complexity increases if one attempts to quantify the importance of a feature on new images…Another reason why noticing Saliency-features does not necessarily facilitate a better understanding of a model is that general-attributes (e.g. colour, contrast) might influence the classification outcome…
Alvarado et al. (2022) How AI Works: Isolated presentation;
Why AI Merits Use: Expressive interaction
Credible Qualitative feedback about the increased ownership of using the system:
p20: “Center in User-Perceived Control and Awareness: …Moreover, participants described an enhanced sense of control and even ‘ownership’ over the recommendation system because TUIs gave them a ‘sense of touching, of power in their hands’.
Emotional Qualitative feedback about the enjoyability of using the system:
p17: “During the post-interview questionnaires, P15 supported these notions: ‘Playing around with them [RD and RP] will very likely improve the experience with the recommendations and, therefore, with the potential enjoyability of the movies/shows.’
Bansal et al. (2019a) How AI Works: Adaptive presentation
Logical Effects of explanations on human mental modal of the error boundary:
p4-5: “…in our controlled experiments we can vary the complexity of the error boundary and measure how it affects the accuracy of humans modeling the true error boundary….In the beginning, the worker makes more mistakes (more red crosses) because the mental model thus far is only partially correct. Eventually, the worker learns the correct model and successfully compensate for the AI (more red checks)…” p6: “…Since error boundary complexity increases with the number of conjunctions, we observe that a more parsimonious error boundary (i.e., a single conjunction) results in a higher team performance.
Bansal et al. (2019b) How AI Works: Adaptive presentation
Logical Effects of explanations on human mental modal of AI predictions:
p2431: “By analogy, we define that an update to an AI component is locally compatible with a user’s mental model if it does not introduce new errors and the user, even after the update, can safely trust the AI’s recommendations.” p2433: “…Figure 3a shows that, as the number of features increases, team performance decreases because it becomes harder to create a mental model…Figure 3b shows that as errors become more stochastic, it becomes harder to create a mental model, deteriorating team performance.
Bansal et al. (2021) How AI Works: Adaptive presentation; Why AI Merits Use: Technical Assurance
Logical Effects of explanations on human agreement on AI predictions:
p9: “Across the three datasets, Adaptive explanations successfully reduced the human’s tendency to blindly trust the AI (i.e., decreased agreement) when it was uncertain and more likely to be incorrect.
Qualitative feedback:
p11(table 3): “…Used AI as a starting references point…Double-checked after they made their own decisions…
Credible Qualitative feedback:
@@
p11(table 3): “…Explicitly acknowledged they used the confidence…Was more likely to accept AI above the threshold
Ben David et al. (2021) How AI Works: Isolated presentation, Protective framing;
Why AI Merits Use: Explanation as recommenders
Logical Effects of explanations on self-reported readiness to adopt and willingness to pay:
p394-5: “On the very first day when the model was introduced (day 3) we see an average of 60% adoption when averaging over all participants…Following the model failure on day 7, adoption plummeted, returning approximately to the initial level with 61% on day 8, this was followed by a slow recovery up until day 10 (66%)…Finally, when participants were told advice will no longer be free, and were asked about willingness to pay, we see an overall decrease in adoption on the final day (77%).
Emotional Algorithmic aversion:
p397: “…the two more elaborate explanations, the Performance-Based (P-value = 0.0445, t = -2.0225) and the Feature-Based (P-value =0.0498, t = -1.9741) prove to be more resistant to the effect of algorithm error…When we explore the adoption recovery…we find that the Performance-Based explanation had a marginally significant (P-value = 0.0633, t = 1.8691) tendency to recover better than the global explanation (12% compared to only 2% for the Feature-Based and the Global Explanation, respectively)…
@@
Bernardo and Seva (2023) How AI Works: Isolated presentation, Explanation affect, Emotional cues
Logical Effects of explanations on self-reported usefulness:
p15: “Example-based explanation increases interestingly surprised, trusting emotions, and perceived usefulness, while feature-and rule-based explanation decreases them. Logic robotic increases fearfully dismayed plus anxiously suspicious emotions and decreases perceived usefulness, while humanized communication functions the opposite. The presence of supplementary information decreases fearfully dismayed emotions, while the absence of it increases the effect.
Credible Effects of explanations on self-reported trust and reliance (see above summary findings)
Emotional Participant AI anxiety and incidental emotion influence user perceived trust of explanations:
p16: “First, when fearfully dismayed emotions were felt, users with high AI anxiety experienced lower perceived trust than those who have low AI anxiety.”
Bove et al. (2022) How AI Works: Mechanistic presentation
Logical Effects of explanations on self-reported system understanding:
p814: “We propose a questionnaire approach…Each item in the questionnaire is a statement, for which users can either answer ”true”, ”false” or ”I don’t know”. We design three types of questions to capture different components of user understanding: (i) explanations’ scope…(ii) explanations’ effects…(iii) explanations’ locality” p816: “Contextualization improves objective understanding…Exploration does not have a significant impact…The interaction of contextualization and exploration does not have a significant impact neither.
Credible Effects of explanations on self-reported satisfaction:
p816: “Contextualization significantly improves users satisfaction…contextualization parameter is greater than the claimed value and conclude that contextualization significantly improves non-expert users’ satisfaction…the exploration parameter is greater than the claimed value and conclude that exploration significantly improves non-expert users’ satisfaction
Cai et al. (2019a) How AI Works: Isolated presentation
Logical Effects of explanations on self-reported system understanding, capability, and benevolence:
p3-4“When the algorithm failed to recognize the drawing, users who received normative explanations felt they had a better understanding of the system…Likewise, when the algorithm did not recognize the drawing, users also rated system capability higher if they had seen a normative explanation…In addition, we found a small but significant main effect of comparative explanations on system benevolence. Users who saw comparative explanations perceived the system to be more benevolent.
Cai et al. (2019b) How AI Works: Mechanistic presentation;
Why AI Merits Use: Explanation as retriever
Logical Effects of explainable interactive tool on self-reported diagnostic utility, future use, and overall preference:
P8(figure 5): “Compared to a conventional interface, SMILY had greater diagnostic utility and required lower effort. Users also indicated having greater trust in SMILY, felt SMILY offered better mental support, and felt they were more likely to use it in clinical practice.
Qualitative feedback:
p10-11: “Tracking the Likelihood of a Decision Hypothesis. During the study, users described ways in which iteratively refining search results helped them track the likelihood of a hypothesis…Generating New Ideas. The refinement tools also provided decision-making support by helping users reflect on their thought process and generate new ideas…Reducing Problem Complexity. Reference images containing a mixture of features were often cognitively difficult to analyze…
Credible Effects of explanations on self-reported trust in the system:
p9: “In sum, users were able to retrieve more diagnostically relevant information with less effort using SMILY, had greater trust in the system, and felt they were more likely to use it in clinical practice…
Cau and Spano (2025) How AI Works: Mechanistic presentation; Why AI Merits Use: Technical assurance
Logical Effects of explanations on accuracy and overreliance:
p1449: “…Consequently, we fail to reject the null hypotheses for H1a and H1b, concluding that the explanation type does not affect users’ accuracy and overreliance on AI…
Self-reported understanding and usefulness of explanations:
p1450-1451: “…15 participants perceived Local Model Explanations (LME) as fast and time-saving support for accepting/rejecting an application…12 participants highlighted the difficulty in understanding why the AI assigned specific weights to attributes and the heavy emphasis on the computer skills attribute…10 participants were overwhelmed by excessive information and suggested improving the visualization of the data distribution module.
Qualitative feedback:
p1451: “…10 participants were overwhelmed by excessive information and suggested improving the visualization of the data distribution module…
Credible Self-reported user confidence and trustworthiness of explanations to use AI:
p1453: “…our exploratory analysis revealed an association between requesting on-demand assistance and lower user confidence. This suggests that accessing data or AI support can undermine confidence, regardless of the type of assistance provided. However, it remains unclear whether explanations reduced confidence or if participants requested assistance due to feeling less confident.
Qualitative feedback:
p1451: “…The Data-Centric Explanations (DCE) were perceived by 18 participants as trustworthy, giving clear guidelines for the job application requirements…
Chen et al. (2023b) How AI Works: Isolated presentation
Logical Effect of explanations on decision-making accuracy and reliance:
p15: “…Across tasks, example-based explanations achieved complementary decision performance whereas feature-based explanations did not…” Qualitative feedback:
P13: “With feature-based explanations, participants applied intuition about features to determine if they agreed or disagreed with the model’s reasoning…example-based explanations can provide additional context and support inductive reasoning
Combined summary:
Example-based explanations led to less disruption of people’s natural intuition about the outcome…Example-based explanations better supported recognizing or even learning new features because the similar examples and their ground truth labels brought in additional context that could promote inductive reasoning…Example-based explanations provided strong signals of unreliability, particularly when the AI system predicts both examples incorrectly.
Díaz-Rodríguez and Pisoni (2020) Why AI Merits Use:
Authority framing, Cultural relatedness
Credible Qualitative feedback:
p319: “…Does embedding of expert/domain knowledge into DL models help explain such models? Can XAI help encode such prior knowledge? Juan Jesus Pleguezuelos, History teacher and podcast author of Art History for entrance exams to University: The challenge I pose is to make others see an historical image only through words. It is clear that this requires an exhaustive description of the masterpiece, but you should also try to make others feel the latent soul in it, and decipher the intention of the author…
Emotional Qualitative feedback:
p319: “…And if you could also convey the emotion that this work is able to cause, it can be that words may be more than enough to make a listener understand an artistic work that he is not seeing in that moment…
Ebermann et al. (2023) How AI Works: Isolated presentation;
Why AI Merits Use: Protective framing
Logical Self-reported evaluation of AI support:
p1815: “This suggests that cognitive misfit reduces the subjective support. However, in contrast to our expectations, the evaluation improved from T2 to T3 in the case of decision situations with cognitive misfit…value on the variable ‘evaluation of AI’s support’ was higher at the third time point of measure…The ‘evaluation of AI’s support’ is indeed lower when cognitive misfit occurs.
Credible Self-reported willingness to use AI:
p1816: “In the case of cognitive misfit resulting from different decisions and corresponding explanations of the AI and the user, users do not have significantly less willingness to use the AI again in comparison to cases where cognitive fit occurs.
Emotional Self-reported mood score:
p1815: “In line with our expectation…where the decision and explanation of the AI is known and cognitive dissonance might occur, that the values of the variable ‘mood’ are significantly lower in the case of participants in decision situations with cognitive misfit…
Ehsan et al. (2021) How AI Works: Isolated presentation; Why AI Merits Use: Social endorsement
Credible Qualitative themes:
Supported Effects…Understanding organizational norms and values, social validation, transitive trust, and temporal relevance…
Ehsan et al. (2024a) How AI Works: Isolated presentation, Numerical association
Logical Self-reported intelligence and understandability:
Participants with AI background also viewed numbers as potentially actionable even when their meaning was unclear…Unable to access their meaning, the non-AI group associated numbers with the presence of a higher, more intelligent expression…both AI and non-AI groups found unanticipated explanatory value in AD’s declarative statements and NR’s numerical representations…
Credible Self-reported confidence, friendliness, and second chance use:
Participants in both groups had unwarranted faith in numbers. However, their extent and reasons for doing so were different…The mere presence of numbers was associated with an algorithmic thinking process in the robot even when the meaning of numbers was unclear…The AI group seems to have followed heuristic reasoning that associates mathematical representations with logic and intelligence…
Evans et al. (2022) How AI Works: Isolated presentation, Graphical synergy;
Why AI Merits: Explanation as recommenders
Logical Self-report on “explanations are understandable” and “explanations are useful for decision-making”:
p286: “Explanation examples are presented in descending order of total positive responses (slightly agree to strongly agree)to the statement: ‘The explanation provides me with valuable information for my work.’ Trust scores were the most widely accepted explanation by this metric, while the explanation receiving the most positive median response was the Counterfactuals (One-axis) example.
Qualitative feedback:
P288-289: “…Two saliency map examples difficult to understand, distracting and/or confusing (p1, 2, 5), to trivially understandable but not at all useful (p6)…The risk for positive confirmation bias with such [saliency] explanations was pointed out: “[It] is hiding a lot of information. But it’s very tempting, because it’s very simple” (p1)
P289: “These [concept] examples were widely considered intuitively understandable and informative about the factors important for the AI output, albeit with limited value ascribed to this information (p1, 2, 4, 6)…Criticisms of the explanation were the lack of precision and/or descriptiveness (p1–3), that having to read text is generally undesirable in the pathologist workflow (p3, 4)
P290: “Pathologists were almost unanimous in expressing that these [counterfactuals] explanations ‘[helped them] understand what the algorithm is looking for’ (p1) All participants drew some concrete conclusions about the factors important to the AI solution, finding it self-explanatory that staining was the most important factor distinguishing positive nuclei from those marked negative, with some identifying other important factors such as shape, size of nuclei, particularly in separating the positively annotated from the unclassified nuclei (in the two-axis example) (p1–3).
Credible Qualitative feedback:
p290: “The labeling of annotations as high- or low-confidence was generally regarded as being intuitively understandable. Aside from identifying the AI solution’s confidence in its annotations, many participants also inferred from this the factors that may have been important to the solution.
Freire et al. (2023) How AI Works: Isolated presentation, Emotional cues
Credible Effects of emoji and explanations on human-AI agreement:
The explainable AI (XAI)-emoji and XAI conditions have higher agreement scores than the emoji and control (no intervention) conditions.
Gaole et al. (2025) How AI Works: Adaptive and Mechanistic presentation
@@
Logical Effects of explanations on user objective feature understanding and appropriate reliance based on human-AI agreement:
p914-915: “We found thatbLLM Agent condition achieved significantly worse objective feature understanding than the Dashboard, CXAI, and ECXAI conditions…In our study, we found that interactive XAI interfaces can have a negative impact of increasing over-reliance on the AI system.
@@
Credible Effects of explanations on self-reported trust of the AI system and user confidence:
p915: “With post-hoc Tukey’s HSD test, we found that participants who received XAI showed significantly higher trust in Understandability/Predictability (i.e., Control ¡ Dashboard, CXAI, ECXAI).”
Emotional Effects of explanations on self-reported engagement of using the system (No direct report in the result)
Kenny et al. (2021) How AI Works: Isolated presentation;
Why AI Merits Use: Explanation as retriever
Logical Effects of explanation on correctness:
p17: “…explanations seem to mainly have an effect when things go wrong, when errors arise and outputs diverge from what a user expects/desires…
Self-reported reasonableness of explanations:
p21: “…this system-level measure of correctness (like the trust measure) shows that the item-level effects for Explanation do not persist into a system-level evaluation of correctness. The Error-Rate’s variable was also found to impact overall correctness and overall reasonableness; see Table 2), showing people are sensitive to changes in error-rates from 4% to 12% and 28%.
Credible Self-reported trust and satisfaction:
p21: “…Specifically, the Explanation-Present group rated Question 1 (“From the explanation, I understand how the program works”), Question 3 (“This explanation of how the program works has sufficient detail”), and Question 5 (“This explanation of how the program works tells me how to use it”) significantly higher than the Explanation-Absent group, all p ¡ .05…
Kim et al. (2024) How AI Works: Mechanistic presentation;
Why AI Merits Use: Explanation as recommenders
Logical Qualitative feedback:
p7-8: “…clinicians found the interfaces helpful in making informed decisions and determining the order of importance of different aspects…To enhance their experience, patients suggested some improvements, including the provision of visual cues to facilitate a more intuitive understanding of their current status. One patient mentioned the following regarding the color-coding scheme…The clinicians expressed a strong desire to identify the types and causes of errors made by their patients…
Kocielnik et al. (2019) How AI Works: Isolated presentation;
Why AI Merits Use: Expectation management